OpenStudy (anonymous):

Solve the following equation by completing the square. 6m2 + 5m -3 = 0

OpenStudy (anonymous):

I could really use some help on this one!

OpenStudy (anonymous):

Firstly, the coefficient of m^2 must be 1.. So, divide by 6 both the sides first... \[m^2 + 5m/6 - 1/2 = 0\] \[(m)^2 + 5m/6 + (5/12)^2 - (5/12)^2 - 1/2 = 0\] \[(m + 5/12)^2 - 97/144 = 0\]

OpenStudy (anonymous):

So how do I find what the "m" term equals?

OpenStudy (anonymous):

To find m, you don't need to do this by completing the square method... Just use the quadratic formula for this...

OpenStudy (anonymous):

Thank you SO much!